OpenStudy (anonymous):

would 12^0 be 0 or 12

OpenStudy (solomonzelman):

\[b^0=1\]

OpenStudy (solomonzelman):

Anything to the power of 0 equals 1.

OpenStudy (solomonzelman):

Do you want to know why?

OpenStudy (anonymous):

yes plz @SolomonZelman

OpenStudy (solomonzelman):

Ok, lets start from dividing and multiplying exponents. \[b^3 \times b^4=b^{3+4}=b^7\]\[v^5-v^2=v^{5-2}=v^3\]those are just example, so are you good with how to divide and multiply exponents?

OpenStudy (solomonzelman):

It should be\[x^5 / v^2= v^{5-2}=v^3\]

OpenStudy (anonymous):

i guess im pretty good

OpenStudy (solomonzelman):

Okl so \[b^0\]can be written as \[b^{a-a}\]OK?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\[b^{a-a}=b^a~ \div~b^a\]so far so good?

OpenStudy (solomonzelman):

@freakenpurplepizza are you good so far? (I'll assume you are busy and can't reply, I'll be back when you do reply if I am online)

OpenStudy (solomonzelman):

Tell me if you get it, till then I'll assume you don't.

OpenStudy (anonymous):

i get it @SolomonZelman

OpenStudy (solomonzelman):

\[So:~~~a^b~\div~a^b=1\]just like any number divided by itself.

OpenStudy (solomonzelman):

Just to summarize what we did. \[b^0=b^{(a-a)}=\frac{b^a}{b^a}=1\]